OpenStudy (anonymous):

intensity of light from bulb is inversely proportional to the square f the distance s from the bulb. if i=90w/m^2 when the distance s=20m, find the itensity at a distance of 10cm/

OpenStudy (anonymous):

formula is .. i=k(1/s) right or not?

OpenStudy (anonymous):

and k=90x20 or k= 1800.. ?

OpenStudy (anonymous):

and.. i sub 2 = 180 w/m^2?

OpenStudy (anonymous):

please correct. :)

OpenStudy (anonymous):

okay walk me through this step by step I think there is something missing here I want to be sure.

OpenStudy (anonymous):

I sub 2= 180? I though it was 1800?

OpenStudy (anonymous):

@godorovg : i sub 2 = 1800 / 10 so.. 180? am i wrong?

OpenStudy (anonymous):

no you are right

OpenStudy (anonymous):

You didn't post i sub 2=1800/10 before that was why I thought it was wrong.

OpenStudy (anonymous):

yeah? oh, ok! Thank you so much! ;D

OpenStudy (anonymous):

no problem I try to teach not give answers I hope I have helped you

OpenStudy (anonymous):

and you are not stupid in math.

OpenStudy (anonymous):

just say because I know I use to feel that way as well, but I can tell you now I am pretty good in math because I learned how math works and I have faith my math skills. Math takes getting your hands dirty and working problems trust me this works..

OpenStudy (anonymous):

it made me sure though :)) yeah, i agree~ haha thanks, again.

OpenStudy (anonymous):

your welcome and if you need help again let me know :)

OpenStudy (anonymous):

gave you a medal